OpenStudy (anonymous):

Using the intermediate value theorem, show that the function g(x)= (2x+1)/(1000x^7-x+1) has at least one point of discontinuity in the interval [-1,0]

OpenStudy (zarkon):

show 1000x^7-x+1 has a root in [-1,0]

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

zarkon is correct

OpenStudy (anonymous):

i dont understand?

OpenStudy (anonymous):

try to plug in (-1,0) to show that 1000^7-x+1 has a root in it

OpenStudy (anonymous):

show that 1000x^7-x+1=0 for x is between [-1,0] so find a value of x between -1 and 0 that makes 1000x^7-x+1=0

OpenStudy (zarkon):

evaluate 1000^7-x+1 at x=-1 ans x=0...then use IVT

OpenStudy (zarkon):

*and

OpenStudy (anonymous):

So only for the denominator? we dont solve for the numerator?

OpenStudy (zarkon):

you have a rational function here ... the only possible cause of a discontinuity is when the denominator is zero